OpenStudy (anonymous):

solve the equation x(x-3)=40

OpenStudy (anonymous):

x(x-3)=40 x^2-3x-40=0 (x-8)(x+5)=0 x=8 or x=-5

OpenStudy (anonymous):

x^2-3x-40=0. By factorization (x-8)(x+5)=0. x=8,-5.

OpenStudy (anonymous):

er x^2-3x-40=0. (x-8)(x+5)=0. x=8,-5.